OpenStudy (anonymous):

Gabrielle is 12 years older than Mikhail. The sum of their ages is 106. What is Mikhail's age?

OpenStudy (anonymous):

41

OpenStudy (anonymous):

47

OpenStudy (anonymous):

g = m + 12 g + m = 106 now sub m + 12 in for g m + 12 + m = 106 (combine like terms) 2m + 12 = 106 2m = 106 - 12 2m = 94 m = 94/2 m = 47

OpenStudy (anonymous):

thank you. !

OpenStudy (anonymous):

will u give medal please